# Involutive Markov categories and the quantum de Finetti theorem

### 15 Dec 2023 v1

@@ 3.2 Almost sure equalities via nullspaces
The implication (ii) $\Rightarrow$ (i) is an immediate application of Theorem 3.18, so we are left to show the converse. For the other direction, by Theorem 3.18 it is enough to show that every $x\in{}_{p}{A}$ can be written as $x=\phi^{\mathrm{op}}(y)-\psi^{\mathrm{op}}(y)$ for some $y\in{}_{p}{C}$ for suitable generalized morphisms $\phi,\psi\colon B\to C$ and suitable $C$. Taking $C=\mathbb{C}^{\{0,1\}}$ and $y=e_{1}$, this is clear from Example 2.29, since it lets us achieve $\phi^{\mathrm{op}}(e_{1})=0$ and $\psi^{\mathrm{op}}(e_{1})=x$. ∎
@@
In the proof of Lemma 3.21, it is not enough to use self-adjoint $\phi$ and $\psi$ only because of Lemma 2.58.
@@ 3.3 Kolmogorov products
    families of generalized morphisms $(\phi_{F}\colon B\to A_{F}\otimes E)$, where $A_{F}$ is shorthand for $\bigotimes_{j\in F}A_{j}$, indexed by finite subsets $F\subseteq J$, such that whenever $F^{\prime}\subseteq F$,
@@
Moreover, this natural bijection must restrict to natural bijections on ${\mathsf{D}}$ and $\operatorname{Gen}({\mathsf{D}})_{\mathrm{det}}$.
@@
Let us consider a picture with countable Kolmogorov products and a morphism $\phi\colon A\to B$. Then we can construct a generalized morphism $\phi^{(\mathbb{N})}\colon A\to B_{\mathbb{N}}$ obtained by the family $\phi^{(n)}$, where $n$ is any nonnegative integer.
@@ 3.4 Representability
By construction, the delta morphisms in $\mathsf{CPU}_{\min/\max}$ correspond to the multiplication maps
@@
In the classical setting, observational representability was introduced in [16, Appendix A.4] based on the notion of observational monad from [34]. The idea behind this concept is that iterated sampling can distinguish any two distinct probability measures. Using 2.40, we can consider the morphisms $\mathsf{samp}^{(n)}_{A}\colon PA\to A^{\otimes n}$ for any $n\geq 1$.
@@
We recall that ${}_{p}{PA}$ is a quotient of the free ***-algebra $F{}_{p}{A}$, as explained in the proof of Proposition 3.36. Therefore, the map $\mathsf{samp}_{A}^{(\mathbb{N}),\mathrm{op}}$ sends by construction an element $x_{1}\odot\dots\odot x_{n}\in{}_{p}{A^{\otimes n}}\subseteq{}_{p}{A_{\mathbb{N}}}$ to the equivalence class of $x_{1}\odot\dots\odot x_{n}$ in ${}_{p}{PA}$. We conclude that $\mathsf{samp}_{A}^{(\mathbb{N}),\mathrm{op}}$ is surjective because every element in ${}_{p}{PA}$ is an equivalence class of finite sums of elements of the form $x_{1}\odot\dots\odot x_{n}$ for some $n$. ∎
@@ 3.5 Classical representability
Let us work out what the $S$-algebra structure on state space objects $SA$ for the pictures $\mathsf{CPU}_{\min/\max}$ amounts to. As we will show, this is exactly the structure which equips the state spaces $\Sigma A$ with the structure of compact convex sets! This is especially important in the foundations of quantum theory, where the convex structure is often considered fundamental due to its simple operational interpretation: a convex combination of states corresponds to a random ‘‘mixing’’ of the states that appear in it. [^15] To see that this was already recognized by von Neumann, note the importance he attributes to discussion of mixtures in his eminent treatise [47]. So although our formal categorical framework does not have any built-in notion of a convex structure, this comes out automatically from classical representability.
To begin, note that under Gelfand duality, there is a functor $C\ell(\mathsf{CPU}_{\min/\max})_{\mathrm{det}}\to\mathsf{CHaus}$ given by completion. In particular, if we restrict the objects to C*-algebras $A={}_{p}{A}$ for simplicity, then this functor is an equivalence. On this subcategory, our monad $S$ corresponds to the Radon monad $P:\mathsf{CHaus}\to\mathsf{CHaus}$, which assigns to every compact Hausdorff space $X$ the space of Radon probability measures $PX$ [22]; this is straightforward to see from the proof of Proposition 3.46 by virtue of the natural isomorphism $\Sigma C(X)\cong PX$ given by the Riesz representation theorem. By a result of Świrszcz [44], the $P$-algebras are exactly the compact convex sets in locally convex spaces, where the structure morphism assigns to every Radon probability measure on such a set its barycenter. In particular, applying the structure morphism to finitely supported probability measures amounts to the formation of convex combinations. So in brief, the $P$-algebras are the compact convex sets. Its free algebras are the probability simplices, which are the closed convex hulls of the Dirac deltas on a compact Hausdorff space $X$ [^16] The precise notion of simplex here is that of Bauer simplex. See [1, Corollary II.4.2] for the theorem which identifies Bauer simplices with the spaces of Radon probability measures on compact Hausdorff spaces.
@@ 3.6 The quantum de Finetti theorem
In this subsection, we consider these questions in the involutive setting. The relevant flavor of representability is classical representability, where the role of distribution objects is now played by state space objects. Based on the quantum de Finetti theorem for states proven as Theorem C.2, we note that these properties apply in our pictures of interest $\mathsf{CPU}_{\min/\max}$.
@@
    in ${\mathsf{D}}$, and this limit is preserved by every $-\otimes\mathrm{id}_{B}$ for every object $B$;
@@
for a unique $\mu\colon C\to QA\otimes B$.
@@
An alternative and categorically more elegant description of Property (ii) of Definition 3.61 is obtained by using a different diagram, given by all the finite tensor powers $A_{F}$ as in Definition 3.24, together with all those deterministic morphisms $A_{F}\to A_{F^{\prime}}$ that are induced from injections $F^{\prime}\hookrightarrow F$. In this way, one can make sense of the universal property of a de Finetti object without requiring the existence of countable Kolmogorov products.
@@
is quite clear: If $\ell_{A}\mathchoice{\,}{\,}{}{}\phi=\ell_{A}\mathchoice{\,}{\,}{}{}\psi$, then also $\ell_{A}^{(\mathbb{N})}\mathchoice{\,}{\,}{}{}\phi=\ell_{A}^{(\mathbb{N})}\mathchoice{\,}{\,}{}{}\psi$ because $\phi$ and $\psi$ are deterministic, and so $\phi=\psi$ by monicity of $\ell_{A}^{(\mathbb{N})}$. To prove surjectivity, consider a morphism $\phi\colon B\to A$. The classicality of $B$ implies that $\phi$ is autocompatible (Remark 2.42(i)), and therefore $\phi^{(\mathbb{N})}\colon B\to A_{\mathbb{N}}$ belongs to ${\mathsf{D}}$ as mentioned in Remark 3.26. By the universal property, we have $\phi^{(\mathbb{N})}=\ell^{(\mathbb{N})}\mathchoice{\,}{\,}{}{}\mu$ for some $\mu\colon B\to QA$. By marginalization, we immediately get $\phi=\ell\mathchoice{\,}{\,}{}{}\mu$, and hence
@@
The pictures $\mathsf{CPU}_{\min/\max}$ are de Finetti representable. In particular, for any morphism $\phi\colon C\to A_{\mathbb{N}}\otimes B$ exchangeable in the first factor, there exists a unique $\mu\colon C\to SA\otimes B$ such that
@@
The submultiplicativity of the norm and the C*-identity on $QA$ also follow easily from (3.6), and therefore $QA$ is a C*-algebra. We define the dense ***-subalgebra as the image of ${}_{p}{A_{\mathbb{N}}}$ under the quotient map, and take $\ell_{A}^{\mathrm{op}}\colon{}_{p}{A}\rightsquigarrow{}_{p}{A_{\mathbb{N}}}\rightsquigarrow{}_{p}{QA}$ to be the composition of the inclusion ${}_{p}{A}\rightsquigarrow{}_{p}{A_{\mathbb{N}}}$ in the first factor and the quotient map ${}_{p}{A_{\mathbb{N}}}\rightsquigarrow{}_{p}{QA}$ itself. Then $\ell_{A}^{\mathrm{op}}$ is unital and contractive, and therefore positive by Proposition A.9. Moreover since $QA$ is commutative, complete positivity follows by Proposition A.13, and also $\ell_{A}$ is trivially autocompatible. Thus Property (i) of Definition 3.61 is satisfied. The monicity of $\ell_{A}\otimes\mathrm{id}_{B}$ for every $B$ required by Property (iii) is clear, since $\ell_{A}^{\mathrm{op}}$ maps ${}_{p}{A_{\mathbb{N}}}$ surjectively onto ${}_{p}{QA}$, and hence also $\ell_{A}\otimes\mathrm{id}_{B}^{\mathrm{op}}$ maps ${}_{p}{A_{\mathbb{N}}}\odot{}_{p}{B}$ surjectively onto ${}_{p}{QA}\odot{}_{p}{B}$.
@@
where $\phi\colon\mathbb{C}\to A_{\mathbb{N}}\otimes B$ ranges over states exchangeable in the first factor. The case $B=\mathbb{C}$ is immediate by definition of $\mathopen{}\mathclose{{}\left\lVert\cdot}\right\rVert_{QA}$. In general, by another application of Bauer’s maximum principle and the quantum de Finetti theorem for states (Theorem C.2),
@@
Combining Lemmas 3.66 and 3.68, we also get the following.
@@ Appendix A From C*-algebras to C*-pairs
Let $(A,{}_{p}{A})$ and $(B,{}_{p}{B})$ be C*-pairs. Then ${}_{p}{A}\odot{}_{p}{B}$ is dense in $A\otimes B$, the C*-algebra obtained by completing $A\odot B$ with respect to any C*-norm (i.e. a submultiplicative norm satisfying the C*-identity).
@@ Appendix C On the quantum de Finetti theorem for states
The arguments of Hulanicki and Phelps on quantum de Finetti theorems [25] are very insightful. In particular, they were able to show that the de Finetti theorem holds in a much more general context than that of C*-algebras [25, Theorem 4.1]. Their method is completely independent of Størmer’s essentially simultaneous paper [43], in which the quantum de Finetti theorem for the minimal tensor product was proven. However, there is a small error in their [25, Theorem 5.4], which implies that their result does not actually apply to the C*-algebra situation: In [25, p. 195], they claim that the space of self-adjoint elements in an infinite tensor power $A_{J}$ is given by the infinite tensor power of the space of self-adjoint elements of $A$. But this is not the case: for example, any element of the form $x\odot y+x^{*}\odot y^{*}$ is self-adjoint even for two non-self-adjoint elements $x$ and $y$. This invalidates their use of [25, Theorem 4.1] in this context.
However, the ideas of Hulanicki and Phelps that enter the proof of their Theorem 4.1 are still applicable. We now show how this results in a quantum de Finetti theorem for both the minimal and the maximal tensor products, which in addition allows for one extra tensor factor in the style of a known criterion for the separability of states in quantum information theory [10, Theorem 1].

### 12 Jan 2024 v2

@@ 3.2 Almost sure equalities via nullspaces
The implication (ii) $\Rightarrow$ (i) is an immediate application of Theorem 3.18, so we are left to show the converse. For the other direction, by Theorem 3.18 it is enough to show that every $x\in{}_{p}{A}$ can be written as $x=\phi^{\mathrm{op}}(y)-\psi^{\mathrm{op}}(y)$ for some $y\in{}_{p}{C}$ for suitable generalized morphisms $\phi,\psi\colon B\to C$ and suitable $C$. For this purpose, we take $C=\mathbb{C}^{\{0,1\}}$ and $y=e_{1}\coloneqq(0,1)$. Taking $C=\mathbb{C}^{\{0,1\}}$ and $y=e_{1}\coloneqq(0,1)$, this is clear from Example 2.29, which lets us achieve $\phi^{\mathrm{op}}(e_{1})=0$ and $\psi^{\mathrm{op}}(e_{1})=x$. ∎
@@
In the proof of Lemma 3.21, $\phi$ and $\psi$ are generally not self-adjoint because of Lemma 2.57.
@@ 3.3 Kolmogorov products
    families of generalized morphisms $(\phi_{F}\colon B\to A_{F}\otimes E)$, where $A_{F}$ is shorthand for $\bigotimes_{j\in F}A_{j}$, indexed by finite subsets $F\subseteq J$, such that whenever $F^{\prime}\subseteq F$, we have
@@
Moreover, this natural bijection is required to restrict to natural bijections on ${\mathsf{D}}$ and $\operatorname{Gen}({\mathsf{D}})_{\mathrm{det}}$.
@@
The notion of Kolmogorov products has already appeared implicitly in the study of the quantum de Finetti theorem by Staton and Summers [42, Definition 2.23 and Theorem 2.24].
@@
Let us consider a picture with countable Kolmogorov products and a morphism $\phi\colon A\to B$. Then we can construct a generalized morphism $\phi^{(\mathbb{N})}\colon A\to B_{\mathbb{N}}$ obtained by the family $(\phi^{(n)})$, where $n$ is any nonnegative integer.
@@ 3.4 Representability
By construction, the sampling morphisms in $\mathsf{CPU}_{\min/\max}$ are simply given by the inclusions $x\mapsto[x]$, while the delta morphisms correspond to the multiplication maps
@@
In the classical setting, observational representability was introduced in [16, Appendix A.4] based on the notion of observational monad from [34]. The idea behind this concept is that iterated sampling can distinguish any two distinct probability measures. Using 2.39, we can consider the morphisms $\mathsf{samp}^{(n)}_{A}\colon PA\to A^{\otimes n}$ for any $n\geq 1$.
@@
We recall that ${}_{p}{PA}$ is a quotient of the free ***-algebra $F{}_{p}{A}$, as explained in the proof of Proposition 3.36. Therefore, the map $\mathsf{samp}_{A}^{(\mathbb{N}),\mathrm{op}}$ sends by construction an element $x_{1}\odot\dots\odot x_{n}\in{}_{p}{A^{\odot n}}\subseteq{}_{p}{A_{\mathbb{N}}}$ to the equivalence class of $x_{1}\odot\dots\odot x_{n}$ in ${}_{p}{PA}$. We conclude that $\mathsf{samp}_{A}^{(\mathbb{N}),\mathrm{op}}$ is surjective because every element in ${}_{p}{PA}$ is an equivalence class of finite sums of elements of the form $x_{1}\odot\dots\odot x_{n}$ for some $n$. ∎
@@ 3.5 Classical representability
Remark 3.46.
Let ${\mathsf{D}}$ be a classically representable picture having countable Kolmogorov products. For any object $A$, the state space object $S(A_{\mathbb{N}})$ is the limit of the diagram $\langle S(A_{F}),S(\pi_{F,F^{\prime}})\rangle$ indexed by finite $F\subseteq\mathbb{N}$ ordered under inclusion. Indeed, $S$ is a right adjoint by Lemma 3.44, and therefore preserves limits. This is a version of [42, Theorem 4.1] formulated in terms of pictures.
@@
Let us work out what the $S$-algebra structure on state space objects $SA$ for the pictures $\mathsf{CPU}_{\min/\max}$ amounts to. As we will show, this is exactly the structure which equips the state spaces $\Sigma A$ with the structure of compact convex sets (cf. [22], and the explicit formulations [42, Theorems 2.18 and 2.19]). This is especially important in the foundations of quantum theory, where the convex structure is often considered fundamental due to its simple operational interpretation: a convex combination of states corresponds to a random ‘‘mixing’’ of the states that appear in it. [^15] To see that this was already recognized by von Neumann, note the importance he attributes to discussion of mixtures in his eminent treatise [47]. So although our formal categorical framework does not have any built-in notion of a convex structure, this comes out automatically from classical representability.
To begin, note that under Gelfand duality, there is a functor $C\ell(\mathsf{CPU}_{\min/\max})_{\mathrm{det}}\to\mathsf{CHaus}$ given by completion. In particular, if we restrict the objects to C*-algebras for simplicity, then this functor is an equivalence. On this subcategory, our monad $S$ corresponds to the Radon monad $P:\mathsf{CHaus}\to\mathsf{CHaus}$, which assigns to every compact Hausdorff space $X$ the space of Radon probability measures $PX$ [22]; this is straightforward to see from the proof of Proposition 3.47 by virtue of the natural isomorphism $\Sigma C(X)\cong PX$ given by the Riesz representation theorem. By a result of Świrszcz [44], the $P$-algebras are exactly the compact convex sets in locally convex spaces, where the structure morphism assigns to every Radon probability measure on such a set its barycenter. In particular, applying the structure morphism to finitely supported probability measures amounts to the formation of convex combinations. So in brief, the $P$-algebras are the compact convex sets. Its free algebras are the probability simplices, which are the closed convex hulls of the Dirac deltas on a compact Hausdorff space $X$ [^16] The precise notion of simplex here is that of Bauer simplex. See [1, Corollary II.4.2] for the theorem which identifies Bauer simplices with the spaces of Radon probability measures on compact Hausdorff spaces.
@@ 3.6 The quantum de Finetti theorem
In this subsection, we consider these questions in the involutive setting. The relevant flavor of representability is classical representability, where the role of distribution objects is now played by state space objects. One theoretical motivation for this is that we would like to factor through an iterated sampling map, and such an operation should be physically implementable, which is not satisfied for general representability (see Remark 3.40). Based on the quantum de Finetti theorem for states proven as Theorem C.2, we note that these properties apply in our pictures of interest $\mathsf{CPU}_{\min/\max}$.
@@
    in ${\mathsf{D}}$, and this limit is preserved by every _$\_\otimes\mathrm{id}_{B}$ for every object $B$;
@@
for a unique morphism $\mu\colon C\to QA\otimes B$.
@@
An alternative and categorically more elegant description of Property (ii) of Definition 3.62 is obtained by using a different diagram, given by all the finite tensor powers $A_{F}$ as in Definition 3.24, together with all those deterministic morphisms $A_{F}\to A_{F^{\prime}}$ that are induced from injections $F^{\prime}\hookrightarrow F$. In this way, one can make sense of the universal property of a de Finetti object without requiring the existence of countable Kolmogorov products. Such an approach is also more consistent with the limit considered by Staton and Summers [42, Lemma 4.2 and Theorem 4.3]. However, Property (ii) seems more natural from an intuitionistic point of view, as it fits better with the narrative of exchangeable morphisms.
@@
is quite clear: If $\ell_{A}\mathchoice{\,}{\,}{}{}\phi=\ell_{A}\mathchoice{\,}{\,}{}{}\psi$, then also $\ell_{A}^{(\mathbb{N})}\mathchoice{\,}{\,}{}{}\phi=\ell_{A}^{(\mathbb{N})}\mathchoice{\,}{\,}{}{}\psi$ because $\phi$ and $\psi$ are deterministic, and so $\phi=\psi$ by monicness of $\ell_{A}^{(\mathbb{N})}$. To prove surjectivity, consider a morphism $\phi\colon B\to A$. The classicality of $B$ implies that $\phi$ is autocompatible (Remark 2.41(i)), and therefore $\phi^{(\mathbb{N})}\colon B\to A_{\mathbb{N}}$ belongs to ${\mathsf{D}}$ as mentioned in Remark 3.26. By the universal property, we have $\phi^{(\mathbb{N})}=\ell^{(\mathbb{N})}\mathchoice{\,}{\,}{}{}\mu$ for some $\mu\colon B\to QA$. By marginalization, we immediately get $\phi=\ell\mathchoice{\,}{\,}{}{}\mu$, and hence
@@
It is worth noting that there is a strong connection between this result and the recent proof of the de Finetti theorem for Markov categories [18, Theorem 4.4]. However, we refrain from discussing this here, as it is beyond the scope of the present paper.
@@
The pictures $\mathsf{CPU}_{\min/\max}$ are de Finetti representable. In particular, for any morphism $\phi\colon C\to A_{\mathbb{N}}\otimes B$ exchangeable in the first factor, there exists a unique morphism $\mu\colon C\to SA\otimes B$ such that
@@
The submultiplicativity of the norm and the C*-identity on $QA$ also follow easily from (3.6), and therefore $QA$ is a C*-algebra. We define the dense ***-subalgebra as the image of ${}_{p}{A_{\mathbb{N}}}$ under the quotient map, and take $\ell_{A}^{\mathrm{op}}\colon{}_{p}{A}\rightsquigarrow{}_{p}{A_{\mathbb{N}}}\rightsquigarrow{}_{p}{QA}$ to be the composition of the inclusion ${}_{p}{A}\rightsquigarrow{}_{p}{A_{\mathbb{N}}}$ in the first factor and the quotient map ${}_{p}{A_{\mathbb{N}}}\rightsquigarrow{}_{p}{QA}$ itself. Then $\ell_{A}^{\mathrm{op}}$ is unital and contractive, and therefore positive by Proposition A.9. Moreover since $QA$ is commutative, complete positivity follows by Proposition A.13, and also $\ell_{A}$ is trivially autocompatible. Thus Property (i) of Definition 3.62 is satisfied. The monicness of $\ell_{A}\otimes\mathrm{id}_{B}$ for every $B$ required by Property (iii) is clear, since $\ell_{A}^{\mathrm{op}}$ maps ${}_{p}{A_{\mathbb{N}}}$ surjectively onto ${}_{p}{QA}$, and hence also $\ell_{A}\otimes\mathrm{id}_{B}^{\mathrm{op}}$ maps ${}_{p}{A_{\mathbb{N}}}\odot{}_{p}{B}$ surjectively onto ${}_{p}{QA}\odot{}_{p}{B}$.
@@
where $\phi\colon\mathbb{C}\to A_{\mathbb{N}}\otimes B$ ranges over states exchangeable in the first factor. The case $B=\mathbb{C}$ is immediate by definition of _$\mathopen{}\mathclose{{}\left\lVert\_}\right\rVert_{QA}$. In general, by another application of Bauer’s maximum principle and the quantum de Finetti theorem for states (Theorem C.2),
@@
Combining Lemmas 3.67 and 3.69, we also get the following.
@@ Appendix A From C*-algebras to C*-pairs
Let $(A,{}_{p}{A})$ and $(B,{}_{p}{B})$ be C*-pairs. Then ${}_{p}{A}\odot{}_{p}{B}$ is dense in $A\otimes B$, the C*-algebra obtained by completing $A\odot B$ with respect to any C*-norm, i.e. a submultiplicative norm satisfying the C*-identity.
@@ Appendix C On the quantum de Finetti theorem for states
The arguments of Hulanicki and Phelps on quantum de Finetti theorems [25] are very insightful. In particular, they were able to show that the de Finetti theorem holds in a much more general context than that of C*-algebras [25, Theorem 4.1]. Their method is completely independent of Størmer’s essentially simultaneous paper [43], in which the quantum de Finetti theorem for the minimal tensor product was proven.
In this appendix, we show how the ideas of Hulanicki and Phelps that enter the proof of their Theorem 4.1 lead to a quantum de Finetti theorem for both the minimal and the maximal tensor products. Moreover, this result allows for one extra tensor factor in the style of a known criterion for the separability of states in quantum information theory [10, Theorem 1].